OpenStudy (anonymous):

Find the value of k so that the line through the given points has the given slope. (-4,2k) and (k, -5); m=-1. PLEASE HELP SO STUCK THANKS

OpenStudy (kirbykirby):

From the slope formula: \(\dfrac{y_2-y_1}{x_2-x_1}\) \[slope =m= \frac{-5-2k}{k-(-4)}=-1 \]

OpenStudy (kirbykirby):

Are you able to solve this type of problem now?

OpenStudy (anonymous):

I am still comfused :)

OpenStudy (kirbykirby):

Do you understand how I applied the slope formula to your problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but what do we do after

OpenStudy (anonymous):

now do we combind

OpenStudy (kirbykirby):

Well you can just solve this like a regular algebra problem. \[\frac{-5-2k}{k-(-4)}=-1\\ ~ \\\frac{-5-2k}{k+4}=-1\] Now cross multiply: \[ -5-2k=-1(k+4)\\ -5-2k=-k=4\] Are you able to solve for k?

OpenStudy (anonymous):

i do not know what you did after -5-2k/k+4=-1. This is the first time learning this.

OpenStudy (anonymous):

cross multiply, but how?

OpenStudy (kirbykirby):

oh if you have a rational equation like... \[ \frac{x+1}{3}=\frac{x+2}{x}\] then what you can do is cross-multiply. That is, the denominator of the 1st fraction multiplies the numerator of the 2nd fraction. And, the denominator of the 2nd fraction multiplies the numerator of the 1st fraction. Or with a picture: